1. (2 points) Find the critical numbers for f(c) = 4x3 - 1522 - 72x + 5
Answer #1

Am f(x) = uni 15 124 +5 For Critical number, f(x) IT नx) = [x) = 4 (4-15-2445) तंत्र बर न 1(x) = 12-x-- 30x -72 = 0 27-5x -